# Learning Chaotic Dynamics with Neuromorphic Network Dynamics

### v1

@@ II.4 Dynamic Reservoir Computing
where $W$ is the static reservoir network weight matrix. In contrast, for neuromorphic reservoirs Eq. (15) is replaced by
@@
where $M(t;\mathbf{r}_{(}t),\mathbf{r}(t-\Delta t),...\mathbf{r}(0);\mathbf{u}(t),\mathbf{u}(t-\Delta t),...,\mathbf{u}(0))$ is a submatrix of $L^{-1}$, where $L$ is the block matrix in Eq. (9). $M$ has size $N\times N_{\rm in}$, such that its elements $M_{ij}=L^{-1}_{ik}$, where $k=j+N$ for $i\in[1,N]$, $j\in[1,N_{\rm in}]$. $M$ is updated at every timestep from the previous $M(t-\Delta t)$ and current input $\mathbf{u}(t)$. Hence, $M$ is dependent on all previous input signals and internal voltage values and encodes all conductance weight changes, via network connectivity and Kirchhoff’s laws.
In conventional reservoirs, all reservoir nodes are linearly combined to generate the output signal. In neuromorphic reservoirs, however, due to the limited number of electrodes used in a physical device, only a subset of all reservoir node voltages, the so called readout nodes, is used to generate the output signal. Another key difference when compared to conventional RCs is the presence of adaptive weights. Although the dynamically evolving conductance values on edges cannot be interpreted as neural network weights, since the neuromorphic network circuitry induces coupled edge–node dynamics instead of a matrix multiplication used in abstract neural networks, they serve a similar role as network weights. Indeed, the most important distinction is the different sources of nonlinearity: nonlinear effects in a conventional reservoir stem from the activation function imposed on its nodes, whereas neuromorphic network nonlinear effects derive from its internal physical dynamics (i.e. electro–ionic transport across memristive edges) and, as we show below, its connectivity.
@@ III Results & Discussion
The results presented below aim to demonstrate how a neuromorphic network works as a physical reservoir and how it is implemented for dynamic RC, using the specific example of autonomous prediction of the Lorenz63 attractor. Both short–term and long–term predictions are presented and analysed, and the influence of external global parameters (input voltage scaling and bias) on forecasting performance is investigated using input–output mappings and a dynamics measure. Nonlinear dynamical features produced by the neuromorphic network are closely examined to determine which features are learned by the auto–regressor used for prediction. How the desirable nonlinear features can be promoted by the set-up of the neuromorphic reservoir is investigated by studying electrode node connectivity and the nanoscale transport dynamics within the memristive edges.
@@ III.1 Lorenz System Prediction
Figure 5: Input–output mapping of a neuromorphic network with $2,000$ nodes and $47,946$ edges. (a) Input signals $\mathbf{r}_{\rm in}(t)$ (blue) constructed from a linear combination of the Lorenz system (cf. Eq. (1)) and delivered to $80$ randomly selected nodes; (b) readout signals $\mathbf{r}_{\rm out}(t)$ (red) from $,91$ 919 randomly selected nodes (distinct from the input nodes). Time duration is the last 10 Lyapunov times of training.
Fig. 5 shows the input–output mapping of a $2,000$–node, $47,946$–edge neuromorphic network used as a reservoir for dynamic RC for Lorenz63 forecasting. Input and output signals are respectively delivered to and read out from $80$ and $,91$ 919 randomly selected nodes. This network and setup was used to autonomously forecast the Lorenz63 signal, with the best example shown in Fig. 6. The forecast time is $t_{f}=9.0$ Lyapunov times, after which the predicted signal deviates but still follows the overall dynamics of the Lorenz attractor. An average forecast time of $2.9\pm 1.2$ Lyapunov times was found over $1,250$ trials, shown in Fig. 7, indicating the network’s general performance in predicting the short-term dynamics of the Lorenz63 system.
Figure 6: Example autonomous prediction of the Lorenz system (5) using a neuromorphic network. (a), (b) and (c) respectively show the individual $\hat{y}_{1}(t)$, $\hat{y}_{2}(t)$ and $\hat{y}_{3}(t)$ values for the predicted signals (orange dashes) and the normalised true Lorenz signals (grey curves). The vertical grey dashed lines indicate the short-term prediction length, defined by Eq. (7), as around $9.0$ Lyapunov times. Data extracted from the same simulation as in Fig. 5.
Figure 7: Histogram of forecast times $t_{f}$, defined by Eq. (7), in units of Lyapunov time for a total of $1,250$ trials using the $2,000$ node neuromorphic network. An average of $2.93$ Lyapunov times is achieved (indicated by the dotted red line), with a standard deviation of $1.21$ Lyapunov times.
@@
Figure 9: Normalised smoothed power spectral density (PSD) of the predicted (red) and true signal (black) of the $y_{3}$ component of the Lorenz system, and the corresponding absolute residual (blue), over a forecast length of $100$ Lyapunov times. Data extracted from the same simulation as in and Fig. 6, during the prediction stage, and smoothed via a Gaussian filter (with a standard deviation of $2$).
@@ III.2 Induced Nonlinear Dynamics
across all time $t_{\ell}$ during training through increments of Lyapunov times (i.e. $t_{\ell+1}-t_{\ell}=\lambda_{\rm max}^{-1}$), normalised by the total number of training timesteps $T$. Here []delimited-[][$\cdot$][ ⋅ ] denote Iverson brackets, where $[P]$ is defined as 1 if statement $P$ is true, and 0 otherwise. The three conductance states are determined by their corresponding $x$ values, with intervals $X_{\rm low}=[0,0.55)$, $X_{\rm med}=[0.55,2/3]$ and $X_{\rm high}=(2/3,1]$ for the low, medium and high conductance states, respectively. Averaging across all $K$ edges in the network, the overall network dynamic measure $\bar{\delta}_{a,b}$ is thus
@@
The output layer weights $W_{\rm out}$ learn nonlinear dynamical features that are most useful for predicting the Lorenz system. Each $i$th column of $W_{\rm out}$ includes the output weighting attributed to the features produced by readout node $i$. Denoting the absolute sum over the $i$th column of $W_{\rm out}$ as $\omega_{i}$, where $\omega_{i}=\sum_{j=1}^{N_{\rm out}}|W_{{\rm out}}^{j,i}|$, a high $\omega_{i}$ implies node $i$ produces more useful features, while low $\omega_{i}$ means that the $i$th feature is suppressed. Fig. 11 shows examples of readout voltages during the last $10$ Lyapunov times of training for three different values of $\alpha$, with readout signals extracted from nodes $i$ with the highest ten $\omega_{i}$ values shown in the top panel (red), and readout signals from the smallest (near zero) ten $\omega_{i}$ values below (magenta), with their corresponding autonomously predicted Lorenz attractor (orange) pictured on the right. It is clear that all cases in Fig. 11 suppress readout signals with high frequency fluctuations, suggesting that such nonlinear features are not useful for learning the Lorenz63 system.
The nonlinear features across different $\alpha$ also show qualitative differences. In Fig. 11(a), the value of $\alpha$ was chosen to match the onset of nonlinearity, at the maxima of the low–med dynamics measure with $\alpha=0.016$V; this is also the case with the largest amount of high frequency fluctuations overall. These fluctuations are not as evident in the two cases (Fig. 11(b) and (c)), with $\alpha=0.2\,{\rm V}$ chosen to match the optimal forecasting time, and with a high value of $\alpha=2.0\,{\rm V}$ which exhibits relatively smoother readouts. These are produced from the persistent dynamics in the formation and decay of conductance paths within the network, shifting the weights within the $M$ matrix gradually at every timestep. The memristive edges evolve across all available conductance states (low through medium to high conductance state, and back), thereby maximising the use of available internal degrees of freedom. This favourable nonlinearity is conducive for good network performance, whereas the high frequency fluctuations represent less useful nonlinearity that is suppressed by $W_{\rm out}$. The corresponding predicted attractor in each case shows that the most useful nonlinear dynamical features are produced away from very low and very high values of $\alpha$, where the predicted attractor becomes unstable.
@@ III.3 Memristor Transport Dynamics
The memristor model, Eq. (11), utilised for this study includes electron tunnelling transport in addition to ballistic transport. As shown in Fig. 3, tunnelling introduces intermediate conductance states that do not exist in the simpler binary model. Knowing that the high frequency fluctuations arise from low to medium conductance state changes (as previously seen in Fig. 11(a)), it is reasonable to question whether they are caused by tunnelling, that is, if the neuromorphic network’s capabilities are limited by the inherent physics of the system.
@@
Fig. 12 reveals, however, that the high frequency fluctuations are even more pronounced in the binary model. The shown examples of both models (simulated with the same parameters) share similar below–average forecast times of $t_{f}\approx 1.6$ Lyapunov times, and their predicted attractors exhibit similar long–term dynamics (up to to 50 Lyapunov times); both successfully reproduce the Lorenz dynamics, only deviating in regions with relatively little training data (at the edges of the attractor). Thus, the high-frequency nonlinear features do not originate from tunnelling dynamics per se, but rather appear to arise as a result of rapid switching of conductance states, fluctuating around the voltage threshold $V_{\rm set}$ for the onset of memristive dynamics. Moreover, the observation that the rapid fluctuations are more prominent in the binary model suggests that the tunnelling model is able to naturally suppress these features, and this may be attributed to the many more degrees of freedom afforded by the intermediate states of tunnelling dynamics (see Fig. A.3 in Appendix). Nevertheless, both models still yield similar average short–term prediction accuracy (see Fig. A.6 in Appendix) due to the suppression of these features via ridge regression in the output layer. See Appendix B for further details on how these high-frequency fluctuations arise from memristive edges in both models.
@@
Figure 13: Examples of (top) memristive edge voltages with and without dynamic activity ($\dot{x}\neq 0$ and $\dot{x}=0$, respectively) and (bottom) their corresponding node voltage readouts, showing that high-frequency nonlinear features may be present even in nodes connected to edges that are not evolving (orange), as compared to a typical nonlinear voltage signal (blue). All simulation parameters are the same as in Fig. 11 with $\alpha=0.2\,{\rm V}$.
@@
The intermediate conductance state changes (low–med, med–high) do not alone explain all fluctuating dynamics in the network; by studying dynamics at the individual memristive edge level, these high frequency fluctuations can also be observed in nodes connected to edges which have no dynamical activity. This is shown in Fig. 13, where high frequency fluctuations can also be observed in edges without any dynamics (i.e. $\dot{x}_{ij}=0$ at edge between node $i$ and node $j$). For a more detailed analysis, see Appendix B. In summary, these high frequency dynamical features are caused in part by the network connectivity distributing nonlinear dynamics amongst all connected parts, effectively acting as a multiplier of nonlinearity, hence allowing network nodes to exhibit nonlinear dynamics even when corresponding edges have no internal dynamical activity themselves.
@@
Overall, nonlinear dynamical features are induced by the complex interactions between memristive edge evolution and network connectivity, resulting in the formation and destruction of conductance paths and feedback loops which cannot be easily inferred unless fully numerically simulated (since a closed-form solution does not exist). From these results, it is expected that proximity of non-active network components to dynamically active components should affect how dynamical features are generated and distributed throughout the network; this is explored next.
@@ III.4 Node Interactions
$m_{\rm nonlin}^{i,j}=\frac{\left\langle{\left|{r}_{\rm out}^{i,j}-{r}_{\rm out}^{L;i,j}\right|^{2}}\right\rangle}{\left\langle{||\mathbf{r}_{\rm out}||^{2}}\right\rangle}.$ (19)
@@
$\bar{m}_{\rm nonlin}=\frac{\left\langle{||\mathbf{r}_{\rm out}-\mathbf{r}_{\rm out}^{L}||^{2}}\right\rangle}{\left\langle{||\mathbf{r}_{\rm out}||^{2}}\right\rangle}.$ (20)
@@
The readout $\mathbf{r}_{\rm out}^{L}$ is obtained from a linear model for which the internal state variables are static with $\dot{\mathbf{x}}=0$ for all memristive junctions, effectively rendering them linear resistors. Consequentially for very small values of $\alpha<0.001\,{\rm V}$, $\bar{m}_{\rm nonlin}=0$ as voltage signals is too small to form any conductance paths (i.e. $\dot{\mathbf{x}}=0$). Likewise for large values of $\alpha>10\,{\rm V}$, the corresponding $\bar{m}_{\rm nonlin}$ are also approximately zero, as all memristive edges are oversaturated with voltage, effectively also resulting in a near linear network with $\dot{\mathbf{x}}\approx 0$ (see Fig. A.4 in Appendix).
@@
Fig. 14 shows heat maps of the time–averaged readout voltages (column 1), along with the low–high dynamics measure (column 2), the sum of the low-med and med-high dynamics measures (column 3), the nonlinearity measure defined in Eq. (19) (column 4), and weight matrix $\omega$ values (column 5), with each pixel value representing a node that respects its relative location on the lattice network. Fig. 14(a) uses uniformly distributed, equidistant input electrode nodes, while Fig. 14(b) and (c) use randomly placed and localised input nodes, respectively. Comparison of these heat maps suggests that the nonlinearity measure is largely correlated with the dynamics measures (especially the low-med and med-high dynamics), and that nonlinearity is anti-correlated with $\omega_{i}$. This anti-correlation implies suppression of highly nonlinear and certain highly dynamical readout nodes by the output layer, which means on a local level some highly nonlinear regions are not useful for the prediction task.
@@
Fig. 14 reveals that nonlinear regions appear mostly in regions which are far away from input electrode nodes. This is more obvious in Fig. 14(c), where by clustering all input nodes together in one region, it becomes clear that the nonlinear parts are the areas away from input nodes. The anti-correlation between $\omega_{i}$ and dynamics measures is also clear in this example, where the highest $\omega_{i}$ weights are mostly in areas near the input nodes. Every input electrode node has its own area of influence, and although this can be affected by other input nodes, on average the further away a readout node is to this input, the less influence this input has on the readout. These far-away nodes dynamically behave in a manner increasingly more different from those of the input electrode nodes.
@@
Note, however, that this anti-correlation is not exact, nor is the correlation between the nonlinearity and dynamics. This is due to the aforementioned network connectivity effects; it is most obvious in the localised input example, where there exists regions with relatively little dynamics but high nonlinearity, and these regions do appear smooth and connected, tending to distribute amongst close neighbouring network components.
@@
The suppressed highly nonlinear nodes also exhibit high frequency fluctuations in their readout signals, which is the unfavourable nonlinearity encountered previously in Fig. 11. Fig. 15 shows the corresponding readout signals of the above three examples in Fig. 14, with the readout signals corresponding with the largest (red) and smallest (magenta) ten $\omega_{i}$ weights both emphasised. As seen previously in Fig. 11, high frequency fluctuations can be observed in Fig. 15 in signals with small $\omega_{i}$ values. As expected, these suppressed nonlinear features are also much more prominent in Fig. 15(c) with the clustered inputs. This implies the unwanted nonlinear readouts arise from nodes distant from the inputs, restricting them to only evolve between the low and medium conductance states as indicated by the low-med dynamic measure, which correlates with the appearance of nonlinear signal that exhibit high frequency fluctuations.
@@
Due to lower presence of unwanted nonlinear features, the examples with random and equidistant inputs both perform better than the clustered input example, as seen from the Lorenz attractor predictions in Fig. 15. This impact of input node placement remains true for the heterogeneous neuromorphic network (see Fig. A.5 in Appendix). Note that with a high network density, randomly selecting nodes leads to an input electrode node placement in which nodes are on average separated by the same number of edges. Even in the sparse lattice network structure there are only minor differences between equidistant and the randomly selected input nodes (cf. Fig. 14(a) and Fig. 14(b)).
@@
To maximise the area of influence of each readout node over time, and hence minimising the occurrence of less useful nonlinear features, the input voltage should remain sufficiently large on average. This can be achieved with the input bias voltages. Instead of sampling values in $\mathbf{b}_{\rm in}$ uniformly from $[-b,b]$, sampling from $[-b,-b/2]\cup[b/2,b]$ ensures that no input node will have low voltages. On average this does have a noticeable effect in the Lorenz prediction task, where for the $500$ node network an average forecast time of $2.12$ Lyapunov times can be observed with the new bias sampling, as opposed to $1.95$ when the bias is sampled form the whole range $[-b,b]$ (cf. Fig. A.7 in Appendix).
@@
In summary, proper input electrode node number and placement allows maximal usage of the given network, by ensuring all readouts are within the area of influence of the inputs such that less useful nonlinear features are minimised. Practically, in a sufficiently dense network, random allocation of input electrode nodes is very similar to evenly distributed input nodes. This motivates our choice to choose the number of input nodes as around $5$% of all available nodes - with roughly evenly spaced input nodes, $5$% is sufficient to activate and continuously sustain network dynamics.

### v2

@@ II.4 Dynamic Reservoir Computing
where $W$ is the static reservoir network weight matrix. In contrast, for the neuromorphic reservoirs considered here, Eq. (15) is replaced by
@@
where $M(t;\mathbf{r}(t),\mathbf{r}(t-\Delta t),...\mathbf{r}(0);\mathbf{u}(t),\mathbf{u}(t-\Delta t),...,\mathbf{u}(0))$ is a submatrix of $L^{-1}$, where $L$ is the block matrix in Eq. (9). $M$ has size $N\times N_{\rm in}$, such that its elements $M_{ij}=L^{-1}_{ik}$, where $k=j+N$ for $i\in[1,N]$, $j\in[1,N_{\rm in}]$. $M$ is updated at every timestep from the previous $M(t-\Delta t)$ and current input $\mathbf{u}(t)$. Hence, $M$ is dependent on all previous input signals and internal voltage values and encodes all conductance weight changes, via network connectivity and Kirchhoff’s laws.
In conventional reservoirs, all reservoir nodes are linearly combined to generate the output signal. However in the physical neuromorphic reservoirs considered here, due to the limited number of electrodes used in a physical device, only a subset of all reservoir node voltages, the so called readout nodes, is used to generate the output signal. Another key difference when compared to conventional RCs is the presence of adaptive weights. Although the dynamically evolving conductance values on edges cannot be interpreted as neural network weights, since the neuromorphic network circuitry induces network dynamics instead of a matrix multiplication used in abstract neural networks, they serve a similar role as network weights. Indeed, the most important distinction is the different sources of nonlinearity: nonlinear effects in a conventional reservoir stem from the mathematical activation function imposed on its nodes, whereas neuromorphic networks in this study derive their nonlinear effects from their internal physical dynamics and heterogeneous connectivity, which influences RC performance [73, 41].
@@ III Results & Discussion
The results presented below aim to demonstrate how a neuromorphic network works as a physical reservoir and how it is implemented for dynamic RC, using the specific example of autonomous prediction of the Lorenz63 attractor. Both short–term and long–term predictions are presented and analysed, and the influence of external global parameters (input voltage scaling and bias) on forecasting performance is investigated using input–output mappings and a dynamics measure. Nonlinear dynamical features produced by the neuromorphic network are closely examined to determine which features are learned by the auto–regressor used for prediction. How the desirable nonlinear features can be promoted by the set–up of the neuromorphic reservoir is investigated by studying electrode node connectivity.
@@ III.1 Lorenz System Prediction
Figure 5: Input–output mapping of a neuromorphic network with $2,000$ nodes and $47,946$ edges. (a) Input signals $\mathbf{r}_{\rm in}(t)$ (blue) constructed from a linear combination of the Lorenz system (cf. Eq. (1)) and delivered to $80$ randomly selected nodes; (b) readout signals $\mathbf{r}_{\rm out}(t)$ (red) from $1,919$ randomly selected nodes (distinct from the input nodes). Time duration is the last 10 Lyapunov times of training.
Fig. 5 shows the input–output mapping of a $2,000$-node, $47,946$-edge neuromorphic network used as a reservoir for dynamic RC for Lorenz63 forecasting. Input and output signals are respectively delivered to and read out from $80$ and $1,919$ randomly selected nodes. This network and setup was used to autonomously forecast the Lorenz63 signal, with the best example shown in Fig. 6. The forecast time is $t_{f}=9.0$ Lyapunov times, after which the predicted signal deviates but still follows the overall dynamics of the Lorenz attractor. An average forecast time of $2.9\pm 1.2$ Lyapunov times was found over $1,250$ trials, shown in Fig. 7, indicating the network’s general performance in predicting the short-term dynamics of the Lorenz63 system.
Figure 6: Example autonomous prediction of the Lorenz63 system (5) using a neuromorphic network. (a), (b) and (c) respectively show the individual $\hat{y}_{1}(t)$, $\hat{y}_{2}(t)$ and $\hat{y}_{3}(t)$ values for the predicted signals (orange dashes) and the normalised true Lorenz signals (grey curves). The vertical grey dashed lines indicate the short-term prediction length, defined by Eq. (7), as around $9.0$ Lyapunov times. The shaded grey region shows the last 5 Lyapunov times of training for a total duration of $\simeq 135$ Lyapunov times, with NRMSE of $0.0012$. Autonomous prediction commences directly after training at $t=0$. Data extracted from the same simulation as in Fig. 5.
Figure 7: Histogram of forecast times $t_{f}$, defined by Eq. (7), in units of Lyapunov time for a total of $1,250$ trials using the $2,000$-node neuromorphic network. An average of $2.93$ Lyapunov times is achieved (indicated by the dotted red line), with a standard deviation of $1.21$ Lyapunov times.
@@
Figure 9: Normalised smoothed power spectral density (PSD) of the predicted (red) and true signal (black) of the $y_{3}$ component of the Lorenz system, and the corresponding absolute residual (blue), over a forecast length of $100$ Lyapunov times. Data extracted from the same simulation as in Fig. 6, during the prediction stage, and smoothed via a Gaussian filter (with a standard deviation of 22).
@@ III.2 Induced Nonlinear Dynamics
across all time $t_{\ell}$ during training through increments of Lyapunov times (i.e. $t_{\ell+1}-t_{\ell}=\lambda_{\rm max}^{-1}$), normalised by the total number of training timesteps $T$. Here $\cdot$] denote Iverson brackets, where $[P]$ is defined as 1 if statement $P$ is true, and 0 otherwise. The three conductance states are determined by their corresponding $x$ values, with intervals $X_{\rm low}=[0,0.55)$, $X_{\rm med}=[0.55,2/3]$ and $X_{\rm high}=(2/3,1]$ for the low, medium and high conductance states, respectively. Averaging across all $K$ edges in the network, the overall network dynamic measure $\bar{\delta}_{a,b}$ is thus
@@
The output layer weights $W_{\rm out}$ learn nonlinear dynamical features that are most useful for predicting the Lorenz system. Each $i$th column of $W_{\rm out}$ includes the output weighting attributed to the features produced by readout node $i$. Denoting the absolute sum over the $i$th column of $W_{\rm out}$ as $\omega_{i}$, where $\omega_{i}=\sum_{j=1}^{N_{\rm out}}|W_{{\rm out}}^{j,i}|$, a high $\omega_{i}$ implies node $i$ produces more useful features, while low $\omega_{i}$ means that the $i$th feature is suppressed. Fig. 11 shows examples of readout voltages during the last $10$ Lyapunov times of training for three different values of $\alpha$, with readout signals extracted from nodes $i$ with the highest ten $\omega_{i}$ values shown in the top panel (red), and readout signals from the smallest (near zero) ten $\omega_{i}$ values below (magenta), with their corresponding autonomously predicted Lorenz attractor (orange) pictured on the right. It is clear that all cases in Fig. 11 suppress readout signals with high–frequency fluctuations, suggesting that such nonlinear features are not useful for learning the Lorenz63 system.
The nonlinear features across different $\alpha$ also show qualitative differences. In Fig. 11(a), the value of $\alpha$ was chosen to match the onset of nonlinearity, at the maxima of the low–med dynamics measure with $\alpha=0.016$V; this is also the case with the largest amount of high–frequency fluctuations overall. These fluctuations are not as evident in the two cases (Fig. 11(b) and (c)), with $\alpha=0.2\,{\rm V}$ chosen to match the optimal forecasting time, and with a high value of $\alpha=2.0\,{\rm V}$ which exhibits relatively smoother readouts. These are produced from the persistent dynamics in the formation and decay of conductance paths within the network, shifting the weights within the $M$ matrix gradually at every timestep. The memristive edges evolve across all available conductance states (low through medium to high conductance state, and back), thereby maximising the use of available internal degrees of freedom. This favourable nonlinearity is conducive for good network performance, whereas the high–frequency fluctuations represent less useful nonlinearity that is suppressed by $W_{\rm out}$. The corresponding predicted attractor in each case shows that the most useful nonlinear dynamical features are produced away from very low and very high values of $\alpha$, where the predicted attractor becomes unstable.
@@ III.3 Memristor Transport Dynamics
The memristor model, Eq. (11), utilised for this study includes electron tunnelling transport in addition to ballistic transport. As shown in Fig. 3, tunnelling introduces intermediate conductance states that do not exist in the simpler binary model. Knowing that the high–frequency fluctuations arise from low to medium conductance state changes (as previously seen in Fig. 11(a)), it is reasonable to question whether they are caused by tunnelling, that is, if the neuromorphic network’s capabilities are limited by the inherent physics of the system.
@@
Fig. 12 reveals, however, that the high–frequency fluctuations are even more pronounced in the binary model. The shown examples of both models (simulated with the same parameters) share similar below–average forecast times of $t_{f}\approx 1.6$ Lyapunov times, and their predicted attractors exhibit similar long–term dynamics (up to to 50 Lyapunov times); both successfully reproduce the Lorenz dynamics, only deviating in regions with relatively little training data (at the edges of the attractor). Thus, the high–frequency nonlinear features do not originate from tunnelling dynamics per se, but rather appear to arise as a result of rapid switching of conductance states, fluctuating around the voltage threshold $V_{\rm set}$ for the onset of memristive dynamics. Moreover, the observation that the rapid fluctuations are more prominent in the binary model suggests that the tunnelling model is able to naturally suppress these features, and this may be attributed to the many more degrees of freedom afforded by the intermediate states of tunnelling dynamics (see Fig. 19 in Appendix). Nevertheless, both models still yield similar average short–term prediction accuracy (see Fig. 23 in Appendix) due to the suppression of these features via ridge regression in the output layer. See Appendix B for further details on how these high–frequency fluctuations arise from memristive edges in both models.
@@
Figure 13: Examples of (top) memristive edge voltages with and without dynamic activity ($\dot{x}\neq 0$ and $\dot{x}=0$, respectively) and (bottom) their corresponding node voltage readouts, showing that high–frequency nonlinear features may be present even in nodes connected to edges that are not evolving (orange), as compared to a typical nonlinear voltage signal (blue). All simulation parameters are the same as in Fig. 11 with $\alpha=0.2\,{\rm V}$.
@@
The intermediate conductance state changes (low–med, med–high) do not alone explain all fluctuating dynamics in the network; by studying dynamics at the individual memristive edge level, these high–frequency fluctuations can also be observed in nodes connected to edges which have no dynamical activity. This is shown in Fig. 13, where high–frequency fluctuations can also be observed in edges without any dynamics (i.e. $\dot{x}_{ij}=0$ at edge between node $i$ and node $j$). For a more detailed analysis, see Appendix B. In summary, these high–frequency dynamical features are caused in part by the network connectivity distributing nonlinear dynamics amongst all connected parts, effectively acting as a multiplier of nonlinearity, hence allowing network nodes to exhibit nonlinear dynamics even when corresponding edges have no internal dynamical activity themselves.
@@
Overall, nonlinear dynamical features are induced by the complex interactions between memristive edge evolution and network connectivity, resulting in the formation and destruction of conductance paths and feedback loops which cannot be easily inferred unless fully numerically simulated (since a closed–form solution does not exist). From these results, it is expected that proximity of non–active network components to dynamically active components should affect how dynamical features are generated and distributed throughout the network; this is explored next.
@@ III.4 Node Interactions
[…]
@@
[…]
@@
The readout $\mathbf{r}_{\rm out}^{L}$ is obtained from a linear model for which the internal state variables are static with $\dot{\mathbf{x}}=0$ for all memristive junctions, effectively rendering them linear resistors. Consequentially for very small values of $\alpha<0.001\,{\rm V}$, $\bar{m}_{\rm nonlin}=0$ as voltage signals is too small to form any conductance paths (i.e. $\dot{\mathbf{x}}=0$). Likewise for large values of $\alpha>10\,{\rm V}$, the corresponding $\bar{m}_{\rm nonlin}$ are also approximately zero, as all memristive edges are oversaturated with voltage, effectively also resulting in a near linear network with $\dot{\mathbf{x}}\approx 0$ (see Fig. 21 in Appendix).
@@
Fig. 14 shows heat maps of the time–averaged readout voltages (column 1), along with the low–high dynamics measure (column 2), the sum of the low–med and med–high dynamics measures (column 3), the nonlinearity measure defined in Eq. (19) (column 4), and weight matrix $\omega$ values (column 5), with each pixel value representing a node that respects its relative location on the lattice network. Fig. 14(a) uses uniformly distributed, equidistant input electrode nodes, while Fig. 14(b) and (c) use randomly placed and localised input nodes, respectively. Comparison of these heat maps suggests that the nonlinearity measure is largely correlated with the dynamics measures (especially the low–med and med–high dynamics), and that nonlinearity is anti–correlated with $\omega_{i}$. This anti–correlation implies suppression of highly nonlinear and certain highly dynamical readout nodes by the output layer, which means on a local level some highly nonlinear regions are not useful for the prediction task.
@@
Fig. 14 reveals that nonlinear regions appear mostly in regions which are far away from input electrode nodes. This is more obvious in Fig. 14(c), where by clustering all input nodes together in one region, it becomes clear that the nonlinear parts are the areas away from input nodes. The anti–correlation between $\omega_{i}$ and dynamics measures is also clear in this example, where the highest $\omega_{i}$ weights are mostly in areas near the input nodes. Every input electrode node has its own area of influence, and although this can be affected by other input nodes, on average the further away a readout node is to this input, the less influence this input has on the readout. These far–away nodes dynamically behave in a manner increasingly more different from those of the input electrode nodes.
@@
Note, however, that this anti–correlation is not exact, nor is the correlation between the nonlinearity and dynamics. This is due to the aforementioned network connectivity effects; it is most obvious in the localised input example, where there exists regions with relatively little dynamics but high nonlinearity, and these regions do appear smooth and connected, tending to distribute amongst close neighbouring network components.
@@
The suppressed highly nonlinear nodes also exhibit high–frequency fluctuations in their readout signals, which is the unfavourable nonlinearity encountered previously in Fig. 11. Fig. 15 shows the corresponding readout signals of the above three examples in Fig. 14, with the readout signals corresponding with the largest (red) and smallest (magenta) ten $\omega_{i}$ weights both emphasised. As seen previously in Fig. 11, high–frequency fluctuations can be observed in Fig. 15 in signals with small $\omega_{i}$ values. As expected, these suppressed nonlinear features are also much more prominent in Fig. 15(c) with the clustered inputs. This implies the unwanted nonlinear readouts arise from nodes distant from the inputs, restricting them to only evolve between the low and medium conductance states as indicated by the low–med dynamic measure, which correlates with the appearance of nonlinear signal that exhibit high–frequency fluctuations.
@@
Due to lower presence of unwanted nonlinear features, the examples with random and equidistant inputs both perform better than the clustered input example, as seen from the Lorenz attractor predictions in Fig. 15. This impact of input node placement remains true for the heterogeneous neuromorphic network (see Fig. 22 in Appendix). Note that with a high network density, randomly selecting nodes leads to an input electrode node placement in which nodes are on average separated by the same number of edges. Even in the sparse lattice network structure there are only minor differences between equidistant and the randomly selected input nodes (cf. Fig. 14(a) and Fig. 14(b)).
@@
To maximise the area of influence of each readout node over time, and hence minimising the occurrence of less useful nonlinear features, the input voltage should remain sufficiently large on average. This can be achieved with the input bias voltages. Instead of sampling values in $\mathbf{b}_{\rm in}$ uniformly from [b,b][-b,b], sampling from $b,-b/2]\cup[b/2,b]$ ensures that no input node will have low voltages. On average this does have a noticeable effect in the Lorenz prediction task, where for the $500$-node network an average forecast time of $2.12$ Lyapunov times can be observed with the new bias sampling, as opposed to $1.95$ when the bias is sampled form the whole range [b,b][-b,b] (cf. Fig. 24 in Appendix).
@@
In summary, proper input electrode node number and placement allows maximal usage of the given network, by ensuring all readouts are within the area of influence of the inputs such that less useful nonlinear features are minimised. Practically, in a sufficiently dense network, random allocation of input electrode nodes is very similar to evenly distributed input nodes. This motivates our choice to choose the number of input nodes as around 55% of all available nodes — with roughly evenly spaced input nodes, 55% is sufficient to activate and continuously sustain network dynamics.
Figure 16: Average forecast time with respect to the number of readout electrode nodes $N_{\rm out}$, for two different sized neuromorphic networks. Shaded region indicates the 25th and 75th percentile, and error bars indicate 1 standard deviation. All data points obtained from 500 trials each.
All numerical simulations performed in this study use $\simeq 90$% of all available nodes as readout nodes, which may limit its hardware implementation. However, as shown in Fig. 16, fewer readouts can still produce good forecast times; for a $2,000$-node network, forecasting saturates above $N_{\rm out}\simeq 1,250$, corresponding to $\simeq 63$%. The readout nodes in Fig. 16 are selected by the size of their $\omega$ weightings (i.e. a portion of the lowest $\omega$ weighted nodes are pruned for small $N_{\rm out}$). Most importantly, Fig. 16 shows that for neuromorphic networks of two different sizes but with the same number of readout nodes, the larger network always outperforms the smaller network: with 500 readouts, the average forecast time is $\simeq 2.4$ Lyapunov times for the larger $2,000$ node network, whereas it cannot exceed 1.8 Lyapunov times for the smaller $500$-node network. Considering the millions of physical nodes and memristive edges in neuromorphic nanowire networks [63, 62], the hardware implementation is expected to vastly outperform the simulations in this study for predicting the Lorenz system. Moreover, experimental [76] and simulation studies [39, 42] indicate that devices can still perform well on physical RC tasks even with a limited number of electrodes. This may be attributed in part to signal integration, since individual electrodes are contacted to many nanowires. Further studies are warranted to determine optimal electrode number and configuration such that they maximise the usage of neuromorphic network dynamics for physical RC.